OpenStudy (anonymous):

i don't get this (6x+2)(2x+8)

OpenStudy (angela210793):

=12x^2+48x+4x+16=12x^2+52x+16

OpenStudy (angela210793):

if this is wht u wanted... 6x*2x+6x*8+2*2x+2*8=12x^2+48x+4x+16=12x^2+52x+16

OpenStudy (anonymous):

that makes no sence

OpenStudy (angela210793):

yes it does

OpenStudy (anonymous):

yea it does